Henrietta Morales is a salaried employee earning $81,600 per year.

Required:
What is Henrietta’s period pay for each of the following pay frequencies? (Round your answers to 2 decimal places.)

Biweekly:

Semimonthly:

Weekly:

Monthly:
Answer #1
Biweekly: 3138.46 =81600/26
Semimonthly: 3400.00 =81600/24
Weekly: 1569.23 =81600/52
Monthly: 6800.00 =81600/12